Harvey Corp has $9mi in total assets, 1$mi in depreciation and amortization expenses. The firm’s basic earning power ratio is 9% and its times interest earned ratio is 3 it has $600,000 in lease payments and $300,000 must go towards principle payments on outstanding loans and L-T debt. Calculate Harvey’s EBITDA coverage ratio.

The correct answer is 2.06 but I don’t know how to solve it
Answer #1

Solution :

The formula for calculating the EBITDA Coverage Ratio is =

( EBITDA + Lease Payments ) / (Lease Payments + Interest payments + Principal Repayments )

As per the information given in the question we have :

Lease Payments = $ 600,000

Principal Repayments on outstanding loans and L-T debt = $ 300,000

Total Assets = $ 9,000,000

Depreciation and Amortization expense = $ 1,000,000

A. As per the information given in the question we have

Earnings Power Ratio = 9 % = 0.09 ; Total Assets = $ 9,000,000

We know that Earnings Power Ratio = ( Earnings before Interest and Taxes i.e., EBIT / Total Assets )

( EBIT / Total Assets ) = 0.09

( EBIT / $ 9,000,000 ) = 0.09

EBIT = 9,000,000 * 0.09 = $ 810,000

Thus EBIT = $ 810,000   - ( A )

B. As per the information given in the question we have

Times Interest Earned Ratio = 3

We know that Times Interest Earned Ratio is calculated using the formula

= EBIT / Interest Expense

( $ 810,000 / Interest Expense ) = 3

Thus Interest Expense = $ 810,000 / 3 = $ 270,000   - ( B )

C. EBITDA = EBIT + Depreciation and Amortization expense

Thus EBITDA = $ 810,000 + $ 1,000,000 = $ 1,810,000   - ( C )

Applying the above values in the formula for EBITDA Coverage Ratio we have :

= ( $ 1,810,000 + $ 600,000 ) / ($ 600,000 + $ 270,000 + + $ 300,000 )

= ( $ 2,410,000 / $ 1,170,000 ) = 2.0598

= 2.06 (when rounded off to two decimal places )

Thus the EBITDA Coverage Ratio of Harvey Corp is = 2.06